the avaerage of two numbers is 11/24. one third of their difference is 1/12. what are these two numbers?
Answer by SadieKhan(31)

Let x and y be the two numbers
As average = x+y/2=11/24
and
(x-y)/3= 1/12
solving 1 and 2
x+y =22/24 = 11/12
x-y = 3/12 = 1/4
from
x+y = 11/12
x= 11/12 - y
putting in the second
11/12-y - y = 1/4
11/12 -2y = 1/4
44/12-8y = 1
44/12 -1 = 8y
8y= (44-12)/12 = 32/12
y= 1/3
as
x+y = 11/12
putting y = 1/3
x+1/3 = 11/12
x = 11/12 - 1/3
x = (11-4)/12
x = 7/12
hence x= 7/12 an y = 1/3